OpenStudy (anonymous):

@esshotwired Which equation is graphed below? A. y = –2x – 2 B. y = 2x C. y = 2x + 2 D. y = 2x – 2 <---

OpenStudy (anonymous):

@esshotwired

OpenStudy (esshotwired):

Yes you are correct again (:

OpenStudy (anonymous):

Thanks! :)

OpenStudy (esshotwired):

You're welcome

OpenStudy (anonymous):

What is the equation of this relation? A. y = 2x + 2 B. y = 2x + 4 <--- C. y = 2x − 2 D. y = 2x − 4 @esshotwired

OpenStudy (esshotwired):

Yep correct again

OpenStudy (esshotwired):

I have to go now so I can't help with anymore for now. Sorry

OpenStudy (anonymous):

@Zale101 @esshotwired @e.mccormick @geerky42 @iambatman @johnweldon1993 @SolomonZelman

OpenStudy (anonymous):

OK! Bye! :) @esshotwired